OpenStudy (anonymous):

Math Analysis: Simplify each expression. (cot^2 x/1 + csc x) + sin x*csc x

OpenStudy (anonymous):

((Cos^2/Sin^2) + (1/Sin)) + Sin*(1/Sin) uhmm typing out cos and sin is annoying, ill use a for sin and b for cos

OpenStudy (anonymous):

can we use a for cos and b for sin? :D

OpenStudy (anonymous):

sure.. hold on

OpenStudy (anonymous):

((a^2/b^2) + (1/b)) + b*(1/b) = ( (a^2/b^2) + (b/b^2) ) + 1 = (a^2+b)/b^2 + 1

OpenStudy (anonymous):

(Cos^2 + Sin)/Sin^2 + 1

OpenStudy (anonymous):

where did you get +1/b from?

OpenStudy (anonymous):

csc = 1/sin

OpenStudy (anonymous):

idk if that can be simplified further. We'll probably end up going around in circles.

OpenStudy (anonymous):

ok so were left with (cos^2 + sin/ sin^2) + 1 so how'd you get csc?

OpenStudy (anonymous):

it was in your question.

OpenStudy (anonymous):

the answer was csc btw at the back of the book but I want to know how you get that

OpenStudy (anonymous):

how would i get that answer im still confused

OpenStudy (anonymous):

sinx*cscx would cancel out of the equation so were left with (cot^2/1 +csc) + 1

OpenStudy (anonymous):

let me do this on paper, im getting confused typing.

OpenStudy (anonymous):

okay lol

OpenStudy (anonymous):

Here you go: \[(Cot^2x)/(1+Cscx)+Sinx*Cscx=\] \[(Cos^2x/Sin^2x)/(1+(1/Sinx))+Sinx*(1/Sinx)=\] \[(Cos^2x/Sin^2x)/(1+(1/Sinx))+1=\] \[(Cos^2x/Sin^2x)/((Sinx+1)/Sinx))+1=\] \[(Cos^2x/Sin^2x)*(Sinx/(Sin+1))+1=\] \[Cos^2x/(Sin^2x+Sinx)+1=\] \[(1-Sin^2x)/(Sin^2x+Sinx)+1=\] \[(1-Sinx)(1+Sinx)/(Sinx(1+Sinx))+1=\] \[(1-Sinx)/Sinx + 1=\] \[(1/Sinx)-(Sinx/Sinx)+1=\] \[(1/Sinx)-1+1=\] \[(1/Sinx)+0=1/Sinx=Cscx\]

OpenStudy (anonymous):

beautiful answer, isn't it :). Pain to type though.

OpenStudy (anonymous):

Still digesting? lol

OpenStudy (anonymous):

yeah its really long o-o

OpenStudy (anonymous):

i dont get (cos^x/sin^2x)*(sinx/sinx + 1) +1 = (cos^2x/sin^2x + sinx) +1 shouldn't it = (cos^2x/sin^2x+1) + 1?

OpenStudy (anonymous):

Where? which line?

OpenStudy (anonymous):

6th line

OpenStudy (anonymous):

Write it out on a piece of paper, you'll see it immediately. I cancelled out Sin on top with Sin^2 on bottom, that left me with a Sin on the bottom which I distributed into the second fractions denominator.

OpenStudy (anonymous):

but where did the 1 on the bottom go?

OpenStudy (anonymous):

so the 1 became a sinx?

OpenStudy (anonymous):

Yes, can't u see i have them under a common denominator? Pay attention to the parenthesis, they are important.

OpenStudy (anonymous):

i still cant understand can you explain how the 1 turned into a sin x I get the rest its just that part. :\

OpenStudy (anonymous):